# From Propeller Damage Estimation and Adaptation to Fault Tolerant Control: Enhancing Quadrotor Resilience

### v1

@@ IV-C Fault-Tolerant Transition
We study the rise-time and effectiveness of our transitioning mechanism by artificially injecting a fault mid-flight. This fault is produced by corrupting a motor input to spin slower. In simulation, we test this mechanism with a $30\%$ and $80\%$ fault. For real-world experiments we inject a $50\%$ fault. Fig. 8 shows the estimation results of propeller damage. When a $30\%$ damage is injected, the fault-tolerant control is not activated and only minor disturbances are observed. However, when an $80\%$ damage fault occurs, the damage estimation of the affected propeller rises to $50\%$ in $0.3\leavevmode\nobreak\ $\mathrm{s}$$, triggering the transition to fault-tolerant control. When the fault-tolerant controller is activated (i.e., $80\%$ damage case), we turn off the estimate and simply extended the last value. For the real-world experiment in Fig. 8, we see a slower response time because the real-world has more noise and the low pass filter requires an increased strength. Fig. 9 shows the corresponding position tracking performances for all cases. We see that the transition to fault-tolerant control results in minor oscillations (up to $10\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}$$) in the $x$ and $y$ positions for both simulation and real-world, but is stable.
@@ V Discussion
The presented results show the benefit of the proposed adaptive control scheme for damage estimation and compensation. However, L1 adaptation can only correct a disturbance if the actions are within the actuator constraints [21]. Therefore a switching threshold for adaptive to fault-tolerant control should reflect this constraint. For example, our system has a $2.5:1$ thrust to weight ratio which enables flight adaptive control with up to $60\%$ damage without switching to fault-tolerant control. We choose to use a switching threshold of $50\%$ to allow some safety margin. For any quadrotor capable of fault-tolerant flight, we assume a $2:1$ thrust to weight ratio minimum is required. Therefore, a conservative switching threshold of $40\%$ damage can be in general considered. Research has shown that L1 adaptive control is effective on a variety of quadrotors from $27$\mathrm{g}$$ Crazyfly [24], $70$\mathrm{g}$$ Parrot Mambo [15], to a heavy weight $700$\mathrm{g}$$ shown in our work along with theoretical guarantees [21]. Our propeller damage estimation method is based on the steady state adaptive properties of L1 [21], making it equally generalizable to other platforms.
One of the current limitations of our approach is that we can only consider single or dual propeller damages cases. Damaged propellers in a near hovering state will spin faster, and undamaged propellers will spin slower or remain at nominal speeds. A reverse effect in hover (lower RPMs of damaged propellers) is highly unlikely, so using the $5\%$ threshold represents a conservative method to estimate propeller damage. Imagine a single damaged propeller that is spinning clockwise. The damaged propeller must spin faster to match the original thrust. However, this produces less torque than the original undamaged propeller creating a yaw acceleration. In order to counteract this yaw acceleration, the two counterclockwise propellers will spin slower. The opposite undamaged clockwise propeller remains consistent with slight deviations. The same principle holds for the case of two damaged propellers. In the adjacent case, a counterclockwise and clockwise torque both decrease. This results in a cancelling effect where the more damaged propeller’s opposite will slow down a constant amount. In the diagonal case, while a similar thrust can be achieved, the alternative spinning undamaged propellers must reduce their corresponding speeds to counteract the yaw acceleration. It is important to note that these actions are handled automatically by our adaption law eq. (14), and we only use our optimization method for calculating the damage estimate.
@@ VI Conclusion
In this paper, we presented an adaptive inference and control strategy to ensure safe quadrotor flight in case of propeller damage. Our approach incorporates an L1 adaptive control technique in conjunction with a fault-tolerant mode. Furthermore, we propose a method to quantify propeller damage using L1 adaptation. Diagnosing the state of the system is vital for safety critical applications as it can help recognize and resolve faults quickly. The proposed method can estimate propeller damage with a significant degree of accuracy. We experimentally study the limits of the L1 adaptive control scheme in this context and analyze at what level of damage severity it is advisable to transition to the fault-tolerant mode.
Future works will attempt recovery and estimation in the middle of more aggressive maneuvers [25]. We would also explore alternative methods to reduce the delay on our low pass filter in real-world tests. Additionally, we would like to see if the proposed approach can be extended to other thrust curve models, especially for flight envelopes where the quadratic relation is no longer a good approximation.

### v2

@@ IV-C Fault-Tolerant Transition
We study the rise-time and effectiveness of our transitioning mechanism by artificially injecting a fault mid-flight. This fault is produced by corrupting a motor input to spin slower. In our experiments, we tested the reliability of our system given 40 trials by injecting $30\%$, $40\%$, and $50\%$ damages in a real drone. $30\%$ damage injection has a $100\%$ success rate of not transition whereas the $50\%$ case shows a $100\%$ successful transitions. The $40\%$ injection has a $87.5\%$ success rate for not transitioning and a $12.5\%$ false positive transition. Our method did not produce any false negative.
Additionally, we inject an $80\%$ damage fault in a real-world experiment to show that our latency is low enough for practical uses. The damage estimation of the affected propeller rises to $50\%$ in $0.15\leavevmode\nobreak\ $\mathrm{s}$$, triggering the transition to fault-tolerant control. Fig. 8 shows the damage estimation in real-world experiments with $30\%,50\%,$ and $80\%$ damage injections along with transition times when relevant. When the fault-tolerant controller is activated (i.e., $>50\%$ damage cases), we stop the inference and estimation process. Fig. 9 shows the corresponding position tracking performances for the non transitioning $30\%$ case and the transitioning $80\%$ damage case. We see our methodology can autonomously transition and stabilize the drone despite the injection of heavy propeller damage. For visualization purposes, the last values are repeated instead of cut when the inference stops after the fault tolerant control is triggered.
@@ V Discussion
The presented results show the benefit of the proposed adaptive control scheme for damage estimation and compensation. However, L1 adaptation can only correct a disturbance if the actions are within the actuator constraints [24]. Therefore a switching threshold for adaptive to fault-tolerant control should reflect this constraint. Our system has a $2.5:1$ thrust to weight ratio which enables flight adaptive control with up to $60\%$ damage without switching to fault-tolerant control. We choose to use a switching threshold of $50\%$ to allow some safety margin. For any quadrotor capable of fault-tolerant flight, a $2:1$ thrust to weight ratio minimum is required. Therefore, a conservative switching threshold of $40\%$ damage can be considered. Research has shown that L1 adaptive control is effective on a variety of quadrotors from $27\leavevmode\nobreak\ $\mathrm{g}$$ Crazyflie [27], $70\leavevmode\nobreak\ $\mathrm{g}$$ Parrot Mambo [15], to our heavy weight platform $700\leavevmode\nobreak\ $\mathrm{g}$$ along with theoretical guarantees [24]. Our propeller damage estimation method is based on the steady state adaptive properties of L1 [24], making it equally generalizable to other platforms.
One limitation of our approach is that we only consider single or dual propeller damages cases. Damaged propellers in a near hovering state will spin faster, and undamaged propellers will spin slower or remain at nominal speeds. A reverse effect in hover (lower RPMs of damaged propellers) is highly unlikely, so using the $5\%$ threshold represents a conservative method to detect propeller damage. Imagine a single damaged propeller that is spinning clockwise. The damaged propeller must spin faster to match the original thrust. However, this produces less torque than the original undamaged propeller creating a yaw acceleration. In order to counteract this yaw acceleration, the two counterclockwise propellers will spin slower. The opposite undamaged clockwise propeller remains consistent with slight deviations. The same principle holds for the case of two damaged propellers. In the adjacent case, a counterclockwise and clockwise torque both decrease. This results in a cancelling effect where the more damaged propeller’s opposite will slow down a constant amount. In the diagonal case, while a similar thrust can be achieved, the alternative spinning undamaged propellers must reduce their corresponding speeds to counteract the yaw acceleration. These actions are handled by our adaption law eq. (11), and our optimization method is only for damage estimation.
@@ VI Conclusion
In this paper, we presented an adaptive inference and control strategy to ensure safe quadrotor flight in case of propeller damage. Our approach incorporates an L1 adaptive control technique in conjunction with a fault-tolerant mode. We propose a method to quantify propeller damage using L1 adaptation that allows smooth transitioning to fault tolerant control in the case of severe damage. The proposed method can effectively account for all range damages from low to severe levels.
Future works will attempt recovery and estimation during more aggressive maneuvers [28]. We would also like to see if the proposed approach can be extended to other thrust curve models, especially for flight envelopes where the quadratic relation is an inaccurate approximation.